In humans, albinism is caused by a recessive allele (a) while normal pigmentation is caused by the dominant allele (A).

a) If two pigmented parents have an albino child, what must the genotypes of the parents be?

b) What is the probability that if they have another child, that child will be albino?

c) What is the probability that the next two children are albinos?

d) What is the probability that they have one albino and one pigmented child? Be careful.

e) If these parents have dizygotic twins, what is the probability that both twins will have the same pigmentation? Again, be careful.
Answer #1

Ans. a. If two pigmented parents have an albino child then both of the parents would be Heterozygous for the gene and would be the carrier of the gene. So, the genotype of both the parents will be Aa.

b. parent genotype =Aa

Aa x Aa

F1 population will be AA, Aa, Aa, aa

Pigmented child probability =3/4

Albino child probability =1/4

C. Probability that next two child are albinos

Probability of albino child is =1/4

Since, they are independent events the probability will be =1/4 x 1/4 = 1/16

d. Probability of pigmented child =3/4

Probability of albino child = 1/4

Since, they are independent events the probability will be = 3/4 x 1/4 = 3/16.